OpenStudy (anonymous):

The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?

OpenStudy (anonymous):

Perpendicular lines have the negative inverse of the other slope

OpenStudy (anonymous):

y=-4x+4 is our original line our perpendicular line is y=1/4x+b 2= 1/4(-8)+b 2= -2+b b=4

OpenStudy (anonymous):

ok thanks